OpenStudy (anonymous):

Which expression is equivalent to this one? u • 5 + u • v A. u(5 + v) B. 5(u + v) C. v(5 + u) D. u · (5 + u) · v

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

by the distributive property \[u\times 5+u\times v=u(5+v)\]

OpenStudy (anonymous):

Thanks ^-^ could you help me with 3 more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

please deposit \(\$.25\) for the next three minutes

OpenStudy (anonymous):

O.o?

OpenStudy (anonymous):

kidding

OpenStudy (anonymous):

unless of course your computer has a vending machine attached

OpenStudy (danjs):

I always think of that old floyd song, I've got a bike, you can ride it if you like... from your pic satellite. lol

OpenStudy (anonymous):

@satellite73 help me pleaasse

OpenStudy (anonymous):

it's got a basket a bell....

OpenStudy (anonymous):

I'm gonna close this question now

OpenStudy (anonymous):

ok bye